4.5m of air a gauge pressure 46.kPa undergoes an isothermal expansion to where the gauge pressure drops to 13.kPa. Calculate the final volume in m3

1
Expert's answer
2022-03-06T15:17:13-0500

Assuming air to be ideal gas and applying Boyle–Mariotte law, have:


"p_1V_1 = p_2V_2"

where "p_1 = 46kPa, p_2 = 13kPa" are initial and final pressure respectively, and "V_1 = 4.5m^3, V_2" are initial and final volumes. Expressing "V_2", obtain:


"V_2 = \dfrac{p_1V_1}{p_2} = \dfrac{46kPa\cdot 4.5m^3}{13kPa} \approx 16m^3"

Answer. 16m^3.